Juwan and Timi Clarke are planning for retirement. Juwan has a number of retirement related questions he needs help answering. Use your retirement planning knowledge to address the following questions.

a.Juwan would like to withdraw $365,000 from retirement savings when he retires. Assuming he can earn a 6.0 percent annualized rate of return on retirement assets, and that inflation will average 2.0 percent during retirement, how much will he need on his first day of retirement to fund twenty-four yearly payments?

b.What will be the inflation-adjusted rate of return (serial rate) if Juwan changes the rate of return and inflation assumptions to be 12.0 percent and 3.0 percent, respectively?

c.Timi will turn age fifty seven this year. She currently contributes to a 401(k) plan and wants to make an IRA contribution. Is she eligible to use the catch-up provision?

d.Juwan is eligible for a modest defined benefit pension when he retires at age sixty-seven. If Juwan elects a 100 percent joint and survivor annuity will he and Timi receive more or less than a 50 percent joint and survivor annuity? If Juwan elects an annuity payment without a survivor benefit, will he receive more or less than a 50 percent joint and survivor annuity? What must occur to obtain a no survivor annuity?

e.Juwan’s employer uses the following unit benefit formula to determine benefits in the defined benefit plan: 2 percent of final pay for each year of service. If Juwan works for twenty years and has a final year income of $160,000 how much will he receive from the pension on a yearly basis?

f.If Timi is age fifty-five and earns $46,000 per year, what is the maximum in annual additions that can be contributed to her qualified 401(k) account this year?
Answer #1

Answer A) This is a case of growing annuity fund

Pu = C* 1-(1+g) *(1 + r)- r-9

r=6% , g= 2% , n=24, C= 365,000

Pv = 365000* 1- (1 +0.02)24 *(1 +0.06) -24 0,06 -0.02= $ 5500098

Answer b) Inflation adjusted return = 12%-3% = 9%

Answer c)The catch-up contribution provision is allowed for people aged 50 or older for additional contributions .

Answer d)i) Under QJSA (qualified Joint and survivor scheme) ,a 100 percent joint and survivor annuity provide the survivor annuity more than 50% and no more than 100%.

ii) Under QJSA (qualified Joint and survivor scheme) ,a an annuity payment without a survivor benefit annuity provide the survivor annuity can maximum up to  50%,

III ) death of annuity holder.